OpenStudy (anonymous):

What is y-5=-2(x+2) ?

OpenStudy (lgbasallote):

expand by distributing -2 into x+2 first...what do you get?

OpenStudy (anonymous):

y-5=-2(x+2) y-5= -2x-4 +5 +5 y=-2x+1